walking and running. On Saturday morning, Lynn walked
for 2 hours and then ran for 30 minutes. If she ran twice as
fast as she walked and she covered 12 miles altogether,
then how fast did she walk?

Answer by stanbon(57246)

Lynn walked for 2 hours and then ran for 30 minutes.
If she ran twice as fast as she walked and she covered 12 miles altogether,
then how fast did she walk?
-------
Walking DATA:
time = 2 hrs; rate = x mph ; distance = 2x miles
-------------------------------
Running DATA
time = 1/2 hr ; rate = 2x mph ; distance = (1/2)(2x) = x miles
-------------------------------
Equation:
distance + distance = 12 miles
2x + x = 12
3x = 12
x = 4 miles (distance she ran)
----
Walking rate = x = 4 mph
================================
Cheers,
Stan H.
================================